Fermi National Laboratory
 

Accelerator Update for May 30, 2006 - May 31, 2006

 

Tuesday May 30

The day shift began with Booster experts conducting a study, with the Antiproton Source (Pbar) in access, with the Tevatron (TeV) at 980 GeV, and with Recycler stashing protons.

Pbar came out of access at 8:36 AM, but when they tried to turn on a breaker for the Debuncher bend bus (D:IB) power supply wouldn’t close. An expert resolved the problem at 9:06 AM.

The Main Injector (MI) resumed ramping their power supplies at 9:13 AM.

Vacuum technicians accessed the TeV at 11:18 AM to check the separator vacuum. Safety system techs also accessed the TeV to begin tests.

The Cryo Coordinator reported 1:36 PM that the wet engine at TeV sector F4 had something loose. Their investigations found a broken connecting rod. The repair will take about ten hours. This will hold off the TeV ramp.

At 3:30 PM, Recycler experts reported Pelletron problems.

TeV accesses ended at 3:32 PM.

Operators had the Collider Detector and Fermilab (CDF) and D-Zero (D0) interlocked by 5:57 PM.

Safety system tests began at 6 PM.

Operators interlocked NuMI at 6:17 PM. NuMI experts began testing their power supplies.

TeV safety system tests ended at 7:39 PM.

Operations resumed the conditioning of the TeV separators at 7:42 PM.

At 8:06 PM, raccoons again tested the cross-gallery defenses. One raccoon was caught by a patrol and questioned. She claimed that she was disoriented by the high frequency noises and accidentally stumbled into the gallery. The patrol took her picture, paw print, and then escorted her out of the building.

Operations began the TeV turn on at 8:37 PM.

Cryo techs reported at 8:47 PM that the F4 wet engine was ready to go.

D0 reported at 9:57 PM that they were powering their solenoid and magnets.

Wednesday May 31

The midnight shift began with Pbar studies, with the Recycler stashing protons, and with TeV at 980 GeV.

Operations began turning off the TeV at 5:02 AM.

Operations began turning off MI at 5:11 AM.

Operations prepared Pbar for access at 5:47 AM.

The Plans for Today

The plans are to access the TeV and other machines today. The bad magnet in the P3 line will be replaced today. The Run Coordinator expects to see beam in the TeV tomorrow.

Machine Reports

Linac

Okay.

Booster

Experts tuned up their machine. Experts will access today to work on the vertical IPM.

MI

Experts tuned to raise the intensity, but more work is needed.

Antiproton Source

Work continues on the stacktail tank electronics. Experts are investigating a possible Debuncher Bend problem.

TeV

All the power supplies have been on and separator conditioning continues. The B11 separator must be vacuum leak checked. The E11 vertical flying wire has a mechanical problem that needs investigating. TeV experts hope to take beam by 8 AM on Thursday, 6/1/06.

Recycler

Experts continue to conduct proton studies. They will access their tunnel today to check equipment. The Electron Cooling made an electron beam, but then the Pelletron broke. An expert investigation will take about two days.

Switchyard 120

Waiting for a B2 dipole magnet installation.

MiniBooNE

Experts are ready for beam.

NuMI

Experts conducted power supply testing. Electricians will access today to make two cable pulls. The dehumidifier units need to be tested further.

Collider Detector at Fermilab

CDF techs have closed up their plug. They had a power supply fail and a gas flow problem; both will be repaired today. CDF experts hope to have the detector closed up and ready for an interlock by 4 PM today.

D-Zero

Experts continue to check out their detector. D0 techs replaced the North Muon shielding and they have run their magnets. They expect the survey crew to be finished by noon today. They will be ready to be interlocked around 6 PM today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/31/2006    email Fermilab